OpenStudy (anonymous):

convert 11pi/12 to degrees

OpenStudy (anonymous):

A full revolution of the unit circle takes \(2\pi\text{ rad}\), or \(360^\circ\). Set up a ratio: \[\frac{2\pi\text{ rad}}{360^\circ}=\frac{\dfrac{11\pi}{12}\text{ rad}}{x^\circ}\] or simply, \[\frac{2\pi}{360}=\frac{11\pi}{12x}\]Solve for \(x\).

OpenStudy (precal):

another quick way to do this is to remember that pi is 180 degrees

OpenStudy (precal):



OpenStudy (precal):

so sub 180 degrees for pi

OpenStudy (precal):

(11 times 180)/12 1980/12 165 degrees that is how I convert from radians to degrees